# OccFusion: A Straightforward and Effective Multi-Sensor Fusion Framework for 3D Occupancy Prediction

### 13 Mar 2024 v2

@@ IV-E Challenging Scenarios Performance Analysis
Method Backbone Input Modality Resolution mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation InverseMatrixVT3D [14] R101-DCN C $200\times 200\times 16$ 12.68 16.33 6.68 0.00 38.88 0.00 15.36 8.07 1.06 0.00 24.24 33.32 2.64 11.64 10.30 4.28 17.37 OccFusion R101+VoxelNet C+R $200\times 200\times 16$ 15.39 12.93 12.50 0.00 48.28 0.00 18.74 9.12 1.28 0.00 29.39 41.57 2.01 18.53 11.63 13.59 26.73 OccFusion R101+VoxelNet C+L $200\times 200\times 16$ 20.77 14.24 13.84 0.00 47.04 0.00 17.67 13.25 1.17 0.00 27.26 44.47 2.54 17.89 28.72 32.36 51.10 OccFusion R101+VoxelNet C+L+R $200\times 200\times 16$ 21.41 17.36 10.66 0.00 51.52 0.00 18.47 14.84 0.68 0.00 31.52 42.83 2.10 18.17 30.86 29.40 52.69
@@
TABLE III: 3D semantic occupancy prediction results on nuScenes validation night scenario subset. All methods are trained with dense occupancy labels from [7]. Notion of modality: Camera (C), Lidar (L), Radar (R).
IV-F Perception Range Impact On Model Performance
@@
Multi-sensor fusion not only improves the final model’s robustness to different illumination and weather conditions but also extends the model’s perception range. We take the vehicle’s centre as the origin and $R$ as the radius. By adjusting the length of $R$, we study the characteristics of different sensor fusion strategies under different perception ranges in different scenarios. We evaluate each model and different sensor fusion strategies at $R=[20m,25m,30m,35m,40m,45m,50m]$.
@@
In the nighttime scenario, the variation trend of performance, as shown in Figure 4(b), demonstrates that adding radar information significantly enhances the lidar’s ability to perceive objects at extended distances. Notably, the performance variation trend of OccFusion(C+L+R) is considerably smoother than that of OccFusion(C+L). This phenomenon can be attributed to radar data having the longest detection range, with velocity information aiding the model in depth prediction. Additionally, despite radar features being sparse, their information can spread to nearby features from other modalities, thereby enhancing the information of the final merged feature.
In the rainy scenario, the performance variation trend, depicted in Figure 4(c), reveals the limited impact of radar information. Specifically, the variation trend of OccFusion(C+L) only slightly deviates from that of OccFusion(C+L+R) when the radius exceeds 30m. Furthermore, the performance variation trend of InverseMatrixVT3D closely aligns with OccFusion(C+R), which further verifies that radar has a limited impact on extending the model perception range under rainy scenarios.
Figure 5: Performance variation trend. (a) performance variation trend on whole nuScenes validation set, (b) performance variation trend on nuScenes validation night scenario subset, and (c) performance variation trend on nuScenes validation rainy scenario subset. Better viewed when zoomed in.
IV-G Framework Qualitative Analysis
@@
We conducted qualitative analysis by generating visualizations of recent state-of-the-art (SOTA) algorithms and comparing them with the prediction results from our framework with different sensor fusion strategies. The overall visualization result is demonstrated in Figure 6. In Figure 6 upper, we present the prediction results for the daytime scenario; in the middle, we show the prediction results for the rainy scenario; and at the bottom, we present the prediction results for the nighttime scenario. The main discrepancy of each prediction result under each scenario has been highlighted by a red rectangle.
@@
In the daytime scenario, as shown in Figure 6 upper, algorithms that rely solely on surround-view cameras are unable to accurately predict pedestrians at remote distances, either failing to identify them or misestimating their numbers. This issue has been partially resolved by integrating radar information with the camera, which means the rada data helps the model extend its perception range. Furthermore, the addition of lidar information has further enhanced the framework’s ability to model the 3D world, particularly in capturing the geometry and contour of static objects.
@@
In the rainy scenario, as shown in Figure 6 middle, vision-centric algorithms have trouble predicting remote range overlay objects. By merging radar information with cameras, this issue has been mitigated a lot, but models still have trouble predicting the manmade buildings that hang in the sky. This issue is well solved by adding lidar data into the model. This visualization result reveals that three-sensor fusion extends the model perception range and enhances the model’s 3D world structure detail-capturing capability.
@@
In the nighttime scenario, the surround-view cameras are susceptible to illumination changes and perform poorly in dim environments. As a result, it is not surprising that purely vision-centric algorithms yield terrible prediction results in such scenarios, as shown in Figure 6 bottom. Surprisingly, we found that even by adding only radar data, which provides a sparse 3D point cloud, the model significantly improves in predicting static objects such as vegetation and manmade structures. Furthermore, when lidar data is merged, the model’s prediction results improve significantly. It is worth noting that even after merging lidar data, the OccFusion (C+L+R) model fails to classify the nearby sidewalk. This phenomenon can be attributed to the fact that lidar sensors do not provide rich semantic information, and in this particular scenario, the camera’s semantic information is also significantly degraded.
IV-H Framework Efficiency Study
@@
We assess the effectiveness of each sensor fusion strategy implemented in our framework and compare them with other state-of-the-art (SOTA) algorithms. Detailed information on the efficiency of the framework is provided in Table IV. By incorporating more sensor information, our framework becomes more complex and necessitates a greater number of trainable parameters. Consequently, this leads to increased utilization of GPU memory and higher latency during inference.
@@
TABLE IV: Model efficiency comparison of different methods. The experiments are performed on a single A10 using six multi-camera images, lidar, and radar data. For input image resolution, all methods adopt $1600\times 900$. $\downarrow$:the lower, the better.
IV-I Framework Ablation Study
IV-I1 Ablation on multi-scale mechanism
We investigated the influence of multi-level supervision and a multi-scale coarse-to-fine feature refinement structure on the overall performance. The ablation study result is shown in Table V.
@@ IV-I1 Ablation on multi-scale mechanism
TABLE V: Ablation study results on the multi-scale mechanism used in the framework. Multi.Stru: multi-scale corse-to-fine refinement structure, Multi.Sup: multi-level supervision mechanism. $\uparrow$:the higher, the better.
The importance of the multi-scale mechanism in improving the performance of the final model is evident. The use of a multi-level supervision approach allows for the deeper levels to capture more general semantic information. Additionally, the multi-scale cores-to-fine refinement structure facilitates the passing of semantic information from deeper to shallower levels, thereby refining the highest-level feature, which is rich in spatial information. Another key finding from the ablation study is the strong correlation between the multi-scale coarse-to-fine structure and the multi-level supervision mechanism. This structure necessitates supervision signals at each scale to enable the framework to capture relevant semantic information and communicate it to higher levels. In the absence of such supervision signals, this design aspect can impede model convergence and result in suboptimal final performance.
IV-I2 Ablation on Dynamic Fusion 3D/2D
@@
TABLE VI: Ablation study results on the dynamic fusion 3D/2D module used in the framework. BEV: merged BEV feature,SENet2D: senet block 2D part used to fuse multi-modality BEV features, SENet3D: senet block 3D part used to fuse multi-modality 3D feature volumes. $\uparrow$:the higher the better.
This study examines the impact of various sub-modules in dynamic fusion 3D/2D on the performance of the final model. Upon removing the BEV feature, thereby disabling the global-local attention fusion module, the model’s performance decreased by approximately 6.5%. This observation suggests that the BEV feature plays a crucial role as an excitation signal, facilitating interaction with the 3D feature volume to enable rapid convergence and capture of 3D geometric information. By removing the SENet3D/2D block in the dynamic fusion module, our model achieves feature fusion solely through feature channel concatenation, devoid of any feature amplification operation. As crucial features remain unamplified, our model experiences a significant performance drop.
@@ V Conclusion
In this study, we present OccFusion, a novel framework that integrates surround-view cameras, radars, and lidar for the prediction of 3D semantic occupancy. Our framework employs dynamic fusion 3D/2D modules to consolidate features from diverse modalities, generating a comprehensive 3D volume. The fusion strategies examined in this research encompass Camera+Radar, Camera+Lidar, and Camera+Lidar+Radar combinations. Through a comprehensive evaluation on the nuScenes validation set, as well as subsets focusing on night and rainy scenarios, we thoroughly analyze the overall performance of each fusion strategy implemented by OccFusion. Additionally, we investigate the impact of perception range on the performance trends of each sensor fusion strategy across different perception ranges. Our experiments exhibit the effectiveness of the OccFusion framework while preserving the distinct strengths of each sensor.

### 09 May 2024 v4

@@ IV-F Challenging Scenarios Performance Analysis
Method Backbone Input Modality IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation OccFusion R101-DCN C 31.10 18.99 18.55 14.29 22.28 30.02 10.19 15.20 10.03 9.71 13.28 20.98 37.18 23.47 27.74 17.46 10.36 23.13 SurroundOcc [7] R101-DCN C 30.57 19.85 21.40 12.75 25.49 31.31 11.39 12.65 8.94 9.48 14.51 21.52 35.34 25.32 29.89 18.37 14.44 24.78 OccFusion R101+VoxelNet C+R 33.75 20.78 20.14 16.33 26.37 32.39 11.56 17.08 11.14 10.54 13.61 22.42 37.50 22.79 29.50 17.58 17.06 26.49 OccFusion R101+VoxelNet C+L 43.36 26.55 24.95 19.11 34.23 36.07 17.01 21.07 18.87 17.46 21.81 28.73 37.82 24.39 30.80 20.37 28.95 43.12 OccFusion R101+VoxelNet C+L+R 43.50 26.72 25.30 18.71 33.58 36.28 17.76 22.44 20.80 15.89 22.63 28.75 39.28 22.72 30.78 20.15 28.99 43.37
@@
TABLE IV: 3D semantic occupancy prediction results on nuScenes validation rainy scenario subset. All methods are trained with dense occupancy labels from [7]. Notion of modality: Camera (C), Lidar (L), Radar (R).
Method Backbone Input Modality IoU mIoU $\bullet$ barrier $\bullet$ bicycle $\bullet$ bus $\bullet$ car $\bullet$ const. veh. $\bullet$ motorcycle $\bullet$ pedestrian $\bullet$ traffic cone $\bullet$ trailer $\bullet$ truck $\bullet$ drive. surf. $\bullet$ other flat $\bullet$ sidewalk $\bullet$ terrain $\bullet$ manmade $\bullet$ vegetation OccFusion R101-DCN C 24.49 9.99 10.40 12.03 0.00 29.94 0.00 9.92 4.88 0.91 0.00 17.79 29.10 2.37 10.80 9.40 8.68 13.57 SurroundOcc [7] R101-DCN C 24.38 10.80 10.55 14.60 0.00 31.05 0.00 8.26 5.37 0.58 0.00 18.75 30.72 2.74 12.39 11.53 10.52 15.77 OccFusion R101+VoxelNet C+R 27.09 11.13 10.78 12.77 0.00 33.50 0.00 12.72 4.91 0.61 0.00 19.97 29.51 0.94 12.15 10.72 11.81 17.72 OccFusion R101+VoxelNet C+L 41.38 15.26 12.74 13.52 0.00 35.85 0.00 15.33 13.19 0.83 0.00 23.78 32.49 0.92 14.24 20.54 23.57 37.10 OccFusion R101+VoxelNet C+L+R 41.47 15.82 13.27 13.53 0.00 36.41 0.00 19.71 12.16 2.04 0.00 25.90 32.44 0.80 14.30 21.06 24.49 37.00
TABLE V: 3D semantic occupancy prediction results on nuScenes validation night scenario subset. All methods are trained with dense occupancy labels from [7]. Notion of modality: Camera (C), Lidar (L), Radar (R).
IV-G Perception Range Impact On Model Performance
@@
Multi-sensor fusion improves the final model’s robustness to illumination and weather conditions and extends the model’s perception range. We take the vehicle’s centre as the origin and $R$ as the radius. By adjusting the length of $R$, we study the characteristics of different sensor fusion strategies under different perception ranges in different scenarios. We evaluate each model and different sensor fusion strategies at $R=[20m,25m,30m,35m,40m,45m,50m]$.
@@
Figure 5(b) depicts the performance variation trend in the rainy scenario. The performance trend of OccFusion(C) displays a notable discrepancy compared to OccFusion(C+R), and this difference becomes more pronounced as the perception range expands. This phenomenon demonstrates the radar sensor’s enhancement to the purely vision-centric algorithms. Nevertheless, the variation trend of OccFusion(C+L) only slightly deviates from that of OccFusion(C+L+R) when the radius exceeds 30m, which shows that in the presence of integrated lidar data in rainy scenes, as the sensing range expands, the contribution of radar decreases.
In the nighttime scenario, the variation trend of performance is shown in Figure 5(c). It demonstrates that adding radar information significantly enhances the framework’s ability to perceive objects at extended distances. Interestingly, the performance disparity between OccFusion (C+L+R) and OccFusion (C+L) widens as the perception range increases, opposite to the trend observed in the rainy scenario.
Figure 5: Performance variation trend for 3D semantic occupancy prediction task. (a) mIoU performance variation trend on the whole nuScenes validation set, (b) mIoU performance variation trend on the nuScenes validation rainy scenario subset, and (c) mIoU performance variation on the nuScenes validation night scenario subset.Better viewed when zoomed in.
IV-H Framework Qualitative Analysis
@@
We conducted qualitative analysis by generating visualizations of recent SOTA algorithms and comparing them with the prediction results from our framework with different sensor fusion strategies. The overall visualization result is demonstrated in Figure 6. In Figure 6 upper, we present the prediction results for the daytime scenario; in the middle, we show the prediction results for the rainy scenario; and at the bottom, we present the prediction results for the nighttime scenario. The red rectangle highlights the main discrepancy of each prediction result under each scenario.
@@
In the daytime scenario, as shown in Figure 6 upper, algorithms that rely solely on surround-view cameras cannot accurately predict pedestrians at remote distances, either failing to identify them or misestimating their numbers. This issue has been partially resolved by integrating radar information with the camera, which means the radar data helps the model extend its perception range. Furthermore, lidar information has further enhanced the framework’s ability to model the 3D world, particularly in capturing the geometry and contour of static objects.
@@
In the rainy scenario, as shown in Figure 6 middle, vision-centric algorithms have trouble predicting remote range overlay objects. This issue has been mitigated by merging radar information with cameras, but models still have trouble predicting the manmade buildings that hang in the sky. This issue can be solved by adding lidar data to the model. This visualization result reveals that three-sensor fusion extends the model perception range and enhances the model’s 3D world structure detail-capturing capability.
@@
In the nighttime scenario, the surround-view cameras are susceptible to illumination changes and perform poorly in dim environments. As a result, it is not surprising that purely vision-centric algorithms yield terrible prediction results in such scenarios, as shown in Figure 6 bottom. Surprisingly, we found that even by adding only radar data, which provides a sparse 3D point cloud, the model significantly improves predicting static objects such as vegetation and manmade structures. Furthermore, when lidar data is merged, the model’s prediction results improve significantly. It is worth noting that even after merging lidar data, the OccFusion (C+L+R) model fails to classify the nearby sidewalk. This phenomenon can be attributed to the fact that lidar sensors do not provide rich semantic information, and in this particular scenario, the camera’s semantic information is also significantly degraded.
IV-I Framework Training Convergence Speed Study
Figure 7: The convergence speed of the framework during the training phase under various sensor fusion strategies. Notion of modality: Camera (C), Lidar (L), Radar (R).
During the training phase, we observed that different sensor fusion strategies influence the framework’s ultimate performance and notably impact its overall training duration. As illustrated in Figure 7, OccFusion (C), a vision-centric method, required 13 training epochs to reach optimal performance. In contrast, OccFusion (C+R), integrating radar information with cameras, reduced the total training epochs to 9 and achieved approximately a 2% performance enhancement. Combining Lidar information with cameras further reduced the training epochs to 6, resulting in a performance improvement of 6%. This occurrence highlights the benefits of sensor fusion, which enhances the final framework performance and accelerates its convergence speed during training.
IV-J Framework Efficiency Study
@@
We assess the effectiveness of each sensor fusion strategy implemented in our framework and compare them with other SOTA algorithms. Table VI provides detailed information on the framework’s efficiency. By incorporating more sensor information, our framework becomes more complex and necessitates more trainable parameters. Consequently, this leads to increased GPU memory utilisation and higher latency during inference.
@@
TABLE VI: Model efficiency comparison of different methods. The experiments are performed on a single A10 using six multi-camera images, lidar, and radar data. For input image resolution, all methods adopt $1600\times 900$. $\downarrow$:the lower, the better.
IV-K Framework Ablation Study
IV-K1 Ablation on multi-scale mechanism
We investigated the influence of multi-level supervision and a multi-scale coarse-to-fine feature refinement structure on the overall performance. The ablation study result is shown in Table VII.
@@ IV-K1 Ablation on multi-scale mechanism
TABLE VII: Ablation study results on the multi-scale mechanism used in the framework. Multi.Stru: multi-scale corse-to-fine refinement structure, Multi.Sup: multi-level supervision mechanism. $\uparrow$:the higher, the better.
The importance of the multi-scale mechanism in improving the mIoU performance of the final model is evident. The use of a multi-level supervision approach allows for the deeper levels to capture more general semantic information. Additionally, the multi-scale coarse-to-fine refinement structure facilitates the passing of semantic information from deeper to shallower levels, thereby refining the highest-level feature, which is rich in spatial information. Another key finding from the ablation study is the strong correlation between the multi-scale coarse-to-fine structure and the multi-level supervision mechanism. This structure necessitates supervision signals at each scale to enable the framework to capture relevant semantic information and communicate it to higher levels. Without such supervision signals, this design aspect can impede model convergence and result in suboptimal final performance.
IV-K2 Ablation on Dynamic Fusion 3D/2D
@@
TABLE VIII: Ablation study results on the dynamic fusion 3D/2D module used in the framework. BEV: merged BEV feature,SENet2D: senet block 2D part used to fuse multi-modality BEV features, SENet3D: senet block 3D part used to fuse multi-modality 3D feature volumes. $\uparrow$:the higher, the better.
This study examines the impact of various sub-modules in dynamic fusion 3D/2D on the performance of the final model. After removing the BEV feature, disabling the global-local attention fusion module, the model’s mIoU performance decreased by approximately 6.7% . This observation suggests that the BEV feature plays a crucial role as an excitation signal, facilitating interaction with the 3D feature volume to enable rapid convergence and capture of 3D geometric and semantic information. By removing the SENet3D/2D block in the dynamic fusion module, our model achieves feature fusion solely through feature channel concatenation, devoid of any feature amplification operation. Our model experiences a significant mIoU performance drop as important features remain unamplified.
@@ V Conclusion
This study presents OccFusion, a novel framework that integrates surround-view cameras, radars, and lidar to predict 3D semantic occupancy. Our framework employs dynamic fusion 3D/2D modules to consolidate features from diverse modalities, generating a comprehensive 3D volume. The fusion strategies examined in this research encompass Camera+Radar, Camera+Lidar, and Camera+Lidar+Radar combinations. Through a comprehensive evaluation on the nuScenes and SemanticKITTI validation set and subsets focusing on night and rainy scenarios, we observed that even a sparse 3D point cloud from surround-view radars can significantly enhance the vision-centric approach. Moreover, the dense 3D point cloud from lidar further improves 3D occupancy prediction performance. Additionally, we explore how the range of perception influences the performance trends of each sensor fusion strategy across varying perception ranges. Our findings reveal that integrating surround-view radars and lidar can significantly enhance the model’s long-distance sensing capabilities and robustness to adverse weather conditions. Our experiments generally exhibit the OccFusion framework’s effectiveness while preserving each sensor’s distinct strengths.